### 16 Mar 2026 v2

# whose only singularity is a double point

Abstract.
We exhibit planar, rational curves of large degree over ${\mathbb{F}}_{2}$ that have a unique singular point, which has multiplicity 2. In characteristic 0 such curves exist only for degrees up to $6$.
The aim of this note is to construct rational curves $C_{d}\subset{\mathbb{F}}_{2}{\mathbb{P}}^{2}$ of degree $d=2^{n}+2$ or $d=2^{n}$ that have a unique singular point, which has multiplicity 2; see Example 3.
In characteristic 0 such curves exist only for degrees up to $6$, see Lemma 4 and Remark 6.
For $d\geq 8$ the curves $C_{d}$ cannot be lifted to characteristic 0 in a way that preserves the standard resolution-of-singularities blow-up sequence; see Remark 7. This is inconsistent with my understanding of some steps of the program outlined in [Ish25]; see Remark 7. However, these curves do not seem to shed light on whether the set of log canonical thresholds depends on the characteristic, which is one of the main topics of [Ish25]; see Remark 8.
A general, planar, rational curve $B_{d}\subset{\mathbb{P}}^{2}$ of degree $d$ has $\binom{d-1}{2}$ nodes, and it is not easy to obtain examples with few singular points. There seem to be very few curves with only one singular point, which is a cusp. These appear prominently in the Abhyankar-Moh-Suzuki theorem describing $\operatorname{Aut}({\mathbb{C}}^{2})$ [Suz74, AM75]. Those with a single Puiseux pair are classified in [FdBLMHN07].
It is especially restrictive to have only cusps with multiplicity 2 [MS89, GZN00]; these can give surprising examples of real polynomials [JK11].
In positive characteristic, Artin-Schreier–type polynomials have been used to give non-standard embeddings of ${\mathbb{A}}^{1}$ into ${\mathbb{A}}^{2}$. This was most likely observed by Abhyankar; [Sat76] refers to them as ‘known counterexamples.’ The curves $C_{d}$ can be given by Artin-Schreier–type equations as in (2.5), but their properties are easier to see using an explicit parametrization.
Example 1.
Fix a characteristic $p>0$. Let $q$ be a $p$-power, and $g(u)\in{\mathbb{F}}_{p}[u]$ a polynomial. Let $C^{\circ}_{q,g}\subset{\mathbb{A}}^{2}$ be the the image of the map
$\phi_{q,g}:t\mapsto\bigl(t^{q},g(t^{p})+t\bigr)=:(x,y),$
and $C_{q,g}\subset{\mathbb{P}}^{2}$ its closure.
Lemma 2.
For the curves $C_{q,g}$ in Example 1 the following hold.
  1. (2.1)
    $\phi_{q,g}:{\mathbb{A}}^{1}\to{\mathbb{A}}^{2}$ is a closed embedding,
  2. (2.2)
    $C^{\circ}_{q,g}$ is a smooth, rational curve, and
  3. (2.3)
    $C_{q,g}$ has a single point at infinity. If $r:=p\cdot\deg g-q$ is positive, it has multiplicity $r$.
Proof. The derivatve of $\phi_{q,g}$ is $(0,1)$, so it is an immersion. Since $t^{q}=x$, it is also purely inseparable, hence an isomorphism. The equation of $C^{\circ}_{q,g}$ is obtained using
(2.4) $y^{q}=g(t^{qp})+t^{q}=g(x^{p})+x.$
If $r:=p\cdot\deg g-q$ is positive, then the unique point at infinity is (0:1:0)(0{:}1{:}0), and the equation there is
(2.5) $z^{r}=g(x^{p})+xz^{q+r-1}.\qed$
Example 3.
Setting $p=2,q=2^{n}$ and $g(u)=u^{2^{n-1}+1}$, we get
$C_{2^{n},2}:=\bigl(z^{2}y^{2^{n}}=x^{2^{n}+2}+xz^{2^{n}+1}\bigr)\subset{\mathbb{P}}^{2}.$
It is a rational curve of degree $2^{n}+2$. It has a unique singular point, which is a cusp of multiplicity 2.
It is thus of type $A_{m}$, where $m=2^{n}(2^{n}+1)$. The latter is roughly $(\deg C_{2^{n},2})^{2}$.
Note that for $m$ even, there are non-equivalent singularities of type $A_{m}$ in characteristic 2 with the same embedded resolution graphs. In the notation of [KS85, GK90], the normal form of the singularity of $C_{2^{n},2}$ is $z^{2}+zx^{r}+x^{m+1}$ where $m=2^{n}(2^{n}+1)$ and $r=\frac{1}{2}m+2^{n-1}+1$.
Setting $p=2,q=2^{n}$ and $g(u)=u^{2^{n-1}-1}$, we get
$C_{2^{n}}:=\bigl(y^{2^{n}}=z^{2}+z^{2^{n}-1}\bigr)\subset{\mathbb{P}}^{2}.$
It is a rational curve of degree $2^{n}$. It has a unique singular point of type $A_{m}$, where $m=(2^{n}-1)(2^{n}-2)$.
Lemma 4.
[GZN00] Let $B\subset{\mathbb{C}}{\mathbb{P}}^{2}$ be a reduced, (not necessarily rational) curve of degree $2d$ with a singularity of type $A_{m}$. Then $m\leq 3d(d-1)+1$, which is roughly $\frac{3}{4}(\deg B)^{2}$.
Hint of proof. Let $S\to{\mathbb{P}}^{2}$ be the double cover ramified along $B$. It has a surface singularity of type $A_{m}$. Thus the cohomology of the vanishing cycles of the $A_{m}$ surface singularity embeds into $H^{2,2}(S_{t},{\mathbb{C}})$ of a smoothing $S_{t}$ of $S$. ∎
Remark 5.
The upper bound in Lemma 4 is not expected to be sharp for $d>6$. The reducible curves
$D_{2d}:=\bigl((y-x^{d})^{2}=y^{2d}\bigr)$
have a singularity of type $A_{2d^{2}-1}$. This corresponds to $m=\frac{1}{2}(\deg D_{2d})^{2}-1$.
It is not easy to find curves $B$ with an $A_{m}$ singularity such that $m>\frac{1}{2}(\deg B)^{2}$. The largest known examples give asymptotically $m\approx\frac{7}{12}(\deg B)^{2}$ [Ore12, Sec.5].
Remark 6.
For $d\leq 6$ [Yos79] gives degree $d$, rational, plane curves $Y_{d}\subset{\mathbb{C}}{\mathbb{P}}^{2}$ with a single singular point, which has multiplicity 2. The curve $Y_{6}$ has an $A_{19}$ singularity; the largest possible. Simpler equations are given in [ABCRCA02, 5.8].
The curve $C_{4,2}$ has degree $6$ and an $A_{20}$ singularity. It is possible that $Y_{6}$ specializes to $C_{4,2}$. Note that the singularity type changes from $A_{19}$ to $A_{20}$, but they both need 10 blow-ups to resolve. So the embedded resolution of $C_{4,2}$ as in (7.1) might be liftable to an embedded resolution of $Y_{6}$.
Remark 7 (Connection with [Ish25]).
Let $0\in X\subset{\mathbb{A}}^{n}$ be a singularity. Assume for simplicity that it can be resolved by a sequence of blowing up points
(7.1) $(Y_{r},X_{r},p_{r})\to\cdots\to(Y_{1},X_{1},p_{1})\to(Y_{0},X_{0},p_{0})=({\mathbb{A}}^{n},X,0),$
where $X_{i}\subset Y_{i}$ is the birational transform of $X$, and $p_{i+1}$ lies over $p_{i}$. Let $\pi_{i}:Y_{i}\to Y_{0}$ denote the composite.
The discrepancies of the exceptional divisors with respect to the ideal sheaf of $X$ are determined by the numbers $\operatorname{mult}_{p_{i}}\pi_{i}^{-1}(p_{0})$ and $\operatorname{mult}_{p_{i}}X_{i}$.
Thus if we deform the blow-up sequence (7.1) such that these multiplicities stay constant, then the discrepancies of the $\pi_{r}$-exceptional divisors are invariant under this deformation. Call these discrepancy preserving deformations of (7.1).
[Ish25] notes that every exceptional divisor over ${\mathbb{A}}^{n}$ can be obtained as a $\pi_{r}$-exceptional divisor for a (somewhat more general) blow-up sequence as in (7.1).
This is especially useful if $0\in X\subset{\mathbb{A}}^{n}$ is in characteristic $p$, and the general fiber of such a deformation is in characteristic 0. It seems that in many cases there are discrepancy preserving liftings from characteristic $p$ to characteristic 0.
However, the curves in Example 3 show that this way of preserving discrepancies in liftings is not always possible.
To get such examples, let $C_{d}=\bigl(g(x,y,z)=0\bigr)$ be a curve of degree $d$ in characteristic $2$ that has a single $A_{2r}$ singularity. For odd $m\gg 1$ consider
$X_{g}:=\bigl(g(x,y,z)+x^{m}+y^{m}+z^{m}=0\bigr)\subset{\mathbb{A}}^{3}.$
This can be resolved by first blowing up the origin, and then $r$-times the singular point of the curve $C_{d}$.
The resulting blow-up sequence has a discrepancy preserving lifting to characteristic 0 iff the curve $C_{d}$ lifts to a curve $C^{\prime}_{d}$ in characteristic 0 that has an $A_{2r}$ or $A_{2r-1}$ singularity.
The smallest counterexample is $C_{8}$, which has degree $8$ and an $A_{42}$ singularity. A reduced, octic curve $B_{8}\subset{\mathbb{C}}{\mathbb{P}}^{2}$ has at most an $A_{37}$ singularity by Lemma 4.
The next one is $C_{8,2}$, which has degree $10$ and an $A_{72}$ singularity. A reduced, decic curve $B_{10}\subset{\mathbb{C}}{\mathbb{P}}^{2}$ has at most an $A_{61}$ singularity by Lemma 4.
Remark 8 (Log canonical thresholds).
The program of Ishii aims to understand log canonical thresholds in positive characteristic [Ish25].
The log canonical threshold of the curve $C_{2^{n},2}\subset{\mathbb{F}}_{2}{\mathbb{P}}^{2}$ is $\frac{1}{2}+\frac{1}{m+1}$, where $m=2^{2n}+2^{n}$. For $n\geq 2$ there is no curve $B\subset{\mathbb{C}}{\mathbb{P}}^{2}$ with the same degree and same log canonical threshold as $C_{2^{n},2}$.
However, the log canonical threshold of the surface $X(g)\subset{\mathbb{A}}^{3}$ constructed from $C_{2^{n},2}$ is $3/(2^{n}+2)$. It is computed by the exceptional divisor of the first blow-up, and a general singular point with the same multiplicity has the same log canonical threshold. Thus these curves do not seem to give further information about the program in [Ish25].
Note, however, that the above non-liftable examples seem rather exceptional, and possibly their discrepancies can be handled in different ways. The next examples show that liftability may frequently hold.
Example 9 (Quartic surfaces).
I first tried to look for non-liftable examples among K3 surfaces. In characteristic 0 the rank of the Néron-Severi group of a smooth K3 is at most 20, but in characteristic $p$ it is 22 for supersingular ones.
So if $S$ is a supersingular K3 surface with Picard number 1 in characteristic $p$, then there is no lifting to characteristic 0 preserving all the singularity types.
Most supersingular K3 surfaces with Picard number 1 have many singularities, and individually each singularity seems to lift, but there are a few examples where lifting even one singularity is not possible. I thank Shimada and Schütt for explaining these to me.
Working with the lists of [Shi04], we discuss 4 extremal cases, but none are good enough to be used in Remark 7.
(9.1) There is a degree 4 K3 surface $S_{1}$ in characteristic $2$ with a $D_{21}$ singularity. However, $S_{1}$ is not a quartic in ${\mathbb{P}}^{3}$, but a double cover of the quadric cone. Thus $S_{1}$ is naturally an octic surface in ${\mathbb{P}}(1,1,2,4)$. So there is an octic surface in ${\mathbb{P}}(1,1,2,4)$ with a $D_{21}$ singularity in characteristic $2$, but there is no such in char 0.
It seems likely that if $(s,S)$ is any quartic surface in characteristic $p$ with with an ADE singularity at $s$, then there is a lifting to characteristic 0 preserving the singularity type at $s$.
(9.2) There is a K3 surface $S_{2}$ in characteristic $11$ with an $A_{21}$ singularity, but it has degree $22$.
(9.3) There is a quasi-elliptic K3 surface $S_{3}$ in characteristic $2$ with a $D_{20}$ singularity, and a section along which the surface is smooth. The Weierstrass equation realizes it is a surface in the projectivization of ${\mathcal{O}}_{{\mathbb{P}}^{1}}+{\mathcal{O}}_{{\mathbb{P}}^{1}}(8)+{\mathcal{O}}_{{\mathbb{P}}^{1}}(12)$, but there is no similar surface in characteristic 0.
(9.4) There is an elliptic K3 surface $S_{4}$ in characteristic $2$ with $A_{17}+3A_{1}$ singularities, and a section along which the surface is smooth. As before, there is no lifting that keeps all singularities, but there is an elliptic K3 surface in characteristic 0 with $A_{17}+A_{1}$ singularities.
The curves $C_{2^{n},2}$ also lead to supersingular double planes.
10Supersingular double planes.
Over ${\mathbb{F}}_{2}$ consider the surfaces
(10.1) $S_{r}:=\bigl(u^{2}+u(zy^{2r}+x^{2r+1})+xz^{4r+1}+z^{2}y^{4r}+x^{4r+2}=0\bigr)\subset{\mathbb{P}}(1,1,1,2r+1).$
In the $y=1$ chart, the equation is
(10.2) $g:=u^{2}+u(z+x^{2r+1})+xz^{4r+1}+z^{2}+x^{4r+2}=0.$
Thus $g_{u}=z+x^{2r+1},\ g_{z}=u+xz^{4r},\ g_{x}=ux^{2r}+z^{4r+1}.$ The first 2 define the monomial curve $M_{r}$ which is the image of
$\phi_{r}:t\mapsto\bigl(t,t^{2r+1},t^{8r^{2}+4r+1}\bigr),$
and $g_{x}$ vanishes identically along $M_{r}$. Thus
$\begin{array}[]{rcl}{\mathbb{F}}_{2}[x,z,u]/(g,g_{x},g_{z},g_{u})&\cong&{\mathbb{F}}_{2}[t]/\bigl(g(t,t^{2r+1},t^{8r^{2}+4r+1})\bigr)\\ &\cong&{\mathbb{F}}_{2}[t]/\bigl(t^{8r^{2}+6r+2}(t^{8r^{2}+2r}+1)\bigr).\end{array}$
Thus $S_{r}$ has an $A_{8r^{2}+6r+1}$ singularity at $t=0$. For the other singularities, write $2r=qr^{\prime}$ where $q$ is a $2$-power and $r^{\prime}$ is odd. Then we have $A_{q-1}$ singularities when $t$ is a $(4rr^{\prime}+r^{\prime})^{th}$ root of unity. There are no other singular points.
If $r$ is a $2$-power, then the minimal resolution $\bar{S}_{r}\to S_{r}$ contains $8r^{2}+6r+1+(4r+1)(2r-1)=16r^{2}+4r$ exceptional curves.
Over the line $(z=0)$ the equation (10.2) becomes $u^{2}+ux^{2r+1}+x^{4r+2}=0$. Setting $v:=ux^{-2r-1}$ we get $v^{2}+v+1=0$. If $\alpha_{1},\alpha_{2}$ are its roots, then the preimage of $(z=0)$ has 2 components
$B_{i}:=\bigl(u+\alpha_{i}x^{2r+1}=z=0\bigr).$
Thus the Picard number of $\bar{S}_{r}$ is at least $16r^{2}+4r+2$.
This equals the second Betti bumber of a double plane ramified along a smooth curve of degree $4r+2$. Thus the surfaces $\bar{S}_{r}$ are supersingular for every $r=2^{n}$, with Picard number $16r^{2}+4r+2$.
Example 11.
The surface $S_{1}$ is a K3 surface with $A_{15}+5A_{1}$ singularities. Resolving the $A_{15}$ singularity, the birational transforms of the $B_{i}$ become smooth rational curves, and the dual configuration is
$\begin{array}[]{ccccccccccccc}&&&&\bar{B}_{1}&&&&\bar{B}_{2}&&&&\\ &&&&|&&&&|&&&&\\ C_{1}&-&C_{2}&-&C_{3}&-&\cdots&-&C_{13}&-&C_{14}&-&C_{15}\end{array}$
From this we compute that $(B_{1}\cdot B_{2})=\frac{9}{16}$ and $(B_{1}^{2})=(B_{2}^{2})=\frac{7}{16}$.
Contracting all but the curve $C_{8}$ gives a supersingular K3 surface $S^{\prime}_{1}$ with singularities $2E_{8}+5A_{1}$, and Picard number $1$. The image of $C_{8}$ is a rational curve $C^{\prime}_{8}$ with 2 cusps at the $E_{8}$-points, and $(C^{\prime}_{8}\cdot C^{\prime}_{8})=2$. Thus $S^{\prime}_{1}$ is a K3 surface of degree 2. This is the 4th surface on the list of [Shi04] in characteristic 2.
Acknowledgments.
I thank M. Schütt and I. Shimada for numerous e-mails explaining supersingular K3 surfaces, and S. Orevkov and M. Zaidenberg for references. Partial financial support was provided by the Simons Foundation (grant number SFI-MPS-MOV-00006719-02).